A sound wave of frequency 300 Hz travels with speed 340ms-1 along









the negative x-direction. Each point of the medium moves up and down through a









total distance of 5.0 mm. Determine the necessary parameters to represent the









wave mathematically.

1
Expert's answer
2022-04-11T12:07:50-0400

"\lambda=\frac v\nu=1.13~m,"

"k=\frac{2\pi}{\lambda}=5.54~m^{-1},"

"\omega=2\pi\nu=1885~\frac{rad}s,"

"y(x,t)=A\sin(kx-\omega t),"

"y(x,t)=0.0025\sin(5.54x-1885t)."